OpenStudy (anonymous):

What is square of 6x^2y^5

OpenStudy (anonymous):

... you mean: \[\LARGE (6x^2y^5)^2 \] or maybe you meant.: SQUARE "ROOT" \[\LARGE \sqrt{6x^2y^5} \] which one? O_O

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

hint... \[\LARGE \left(x^a\right)^b=x^{a\cdot b}\] I guess you can solve it now :)

OpenStudy (anonymous):

36x^4y^10 ?

OpenStudy (anonymous):

yes.. that's correct. Well done :)